# Privacy-Aware Detection of Fake Identity Documents: Methodology, Benchmark, and Improved Detection Methods (FakeIDet2)

### v1

@@ 2.2 Fake ID Detection Methods and International Challenges
One of the first methods was presented in [22], where the authors trained a CNN via pixel-wise classification using an in-house database, reporting a 2.22% EER. Gonzalez and Tapia proposed in [13] a two-stage system which first used a neural network to evaluate if the submitted ID is real or fake, evaluating digital PAIs of type composite or synthetic. The second network classified the submitted ID between real or fake, evaluating physical PAIs such as print, display or PVC. When concatenating both systems, the Bona-fide Presentation Classification Error Rate at a fixed decision threshold $\tau=0.01$($BPCER_{100}$) is 0.92%.
@@
Evidences that the fake ID detection field is receiving increasing attention in the last years is the organization of international challenges in top conferences. The first challenge was introduced at IEEE IJCB 2024 [33]. No data was given by the organizers, and the submitted fake ID detectors were evaluated over real IDs, which further reduced the model’s performance, since the winner of the challenge achieved an EER of 21.87%. The second edition of this challenge has been hosted at IEEE IJCB 2025666https://sites.google.com/view/ijcb-pad-id-card-2025/, where the organizers proposed two tracks. The first track considered IDs created by the organizers using PVC cards as “real” documents and the rest as attacks, and the second track treated only official IDs as real. Only for the first track, a database was given to the participants. The winner for the first track reported an EER of 11.34%, while for the second track, the best team reported an EER of 6.36%, but the second and third teams reported EER s of 23.87% and 31.94%, respectively. These results showcase this is a challenging task, specially due to the lack of public real data. Recently, the first challenge including digital manipulations for fake ID detection has been organized at IEEE/CVF ICCV 2025777https://deepid-iccv.github.io/#home. This challenge proposed two tracks, one for binary classification, which focused on PAD (i.e., detecting fake and real documents) and one for localization (i.e., providing a mask that showed the tampered sections). The organizers provided a database for all participants, named FantasyID, which contained fake templates of different countries from different regions of the world. The “real” ID data was created by printing high-quality fakes using PVC cards, and taking images from them. After that, those images were digitally altered through face-swapping and text inpainting methods. For the detection track the evaluation was performed using the F1-score with a decision threshold fixed to $\tau=0.5$ over two datasets: the FantasyID evaluation set, which contained PAI s which were not available in the training set, and a private dataset which was provided by the company PXL Vision, with the same type of attacks as in the FantasyID dataset, but using official IDs as real. The final score, denominated Aggregated F1-score, was a pondered mean that gave more importance to the F1-score obtained over the private dataset, hence, benefiting models that generalized better to out-of-distribution data. The winning team achieved an Aggregated F1-score of 0.8, obtaining a F1-score of 0.99 and 0.72 in both the FantasyID dataset and private dataset, respectively. These results, although impressive, show that there is a present issue in domain adaptation between official real IDs and fake samples created in laboratory conditions that are considered “real”. Hence, providing privacy preserving mechanisms that allow to share official real IDs without compromising individual’s personal information is fundamental to further advance in this research field.
2.3 State-of-the-Art: Summary
@@
A clear trend can be observed based on previous sections. Public databases in the literature do not have official real data available, and fake ID detectors trained using private databases with real data do not release neither the code nor the weights to the research community, making impossible to perform fair comparisons among state-of-the-art methods. Moreover, the characteristics of those databases are not always detailed (e.g., acquisition devices or number of unique IDs), which make even more difficult to advance in the field. The present study aims to advance through the following key contributions: i) the release of a novel database, FakeIDet2-db, increasing the number of images from real and fake IDs, ii) the inclusion of a new kind of PAIs, named physical composite, to the proposed database, being the first one that contains this kind of attack, iii) a novel method for privacy-aware fake ID detection based on fusion of representations from individual patches, and iv) a novel standard public benchmark that considers a wide variety of physical and synthetic attacks included in other public databases in the literature, considering IDs with different templates and demographic regions.
@@ 3 Proposed Database: FakeIDet2-db
This section describes our novel FakeIDet2-db. Concretely, Sec. 3.1 and Sec. 3.2 provide all the details of the real and fake ID acquisition, whereas Sec. 3.3 describes the post-processing carried out for the different anonymization and patch size configurations.
@@ 3.1 Acquisition Setup: Real IDs
Figure 3: Main characteristics of our proposed public FakeIDet2-db. *Note ‐ For the Fake IDs - Attacks, the personal information is redacted in this figure to protect the privacy of the subjects. Official Real IDs are based on non-real examples.
Fig. 3 summarizes the main characteristics of our proposed public FakeIDet2-db. Regarding the number of unique IDs, our database comprises 47 different official Spanish IDs. Spanish electronic IDs appearance have changed during the years, which is a key aspect that has also been considered while capturing this database, including three different versions. For the acquisition, we used three different smartphone models (Apple iPhone 15, Xiaomi Mi 9T Pro and Redmi 9C NFC), aiming to cover the range of sensor quality from high-end to low-end, respectively. Their main technical specifications are provided in Table 2 where we can see that both iPhone 15 and Mi 9T Pro have main camera sensors with 48MP while the Redmi 9C NFC only has 13MP. The iPhone camera is the best one in terms of adaptability to low light scenarios, given its aperture of ff/1.6, although Mi 9T Pro is close with an aperture of ff/1.75. However, the Redmi 9C NFC struggles in these kind of low-light scenarios, given that its aperture is greater, ff/$2.$ Regarding image resolution, the iPhone 15 is able to shoot ProRes images up to 8K resolution, but for our acquisition process, we set it to 4K, since both Mi 9T Pro and Redmi 9C NFC can only shoot images at that resolution. Pictures of IDs were taken using a 4:3 aspect ratio. Three different heights were considered with respect to the document’s vertical: 10 cms, 12.5 cms and 15 cms, building a 3D printed adjustable platform. This enabled acquiring IDs precisely, similar to [22], where they captured IDs using a mobile application that provided a template in the camera preview to match where the IDs borders should be when the picture was taken.
In order to provide more variability, we also considered different light conditions in terms of room illumination: i) no-light, with the camera flash activated, ii) dim-light, with images taken with/without the camera flash activated, and iii) bright-light, with images taken with/without the camera flash activated.
Therefore, given the aforementioned variability factors, for a unique ID, we obtained 45 images (i.e., 3 smartphones $\times$ 3 heights $\times$ 5 light-conditions) covering different conditions, yielding a total of 2,115 (i.e., 47 $\times$ 45) images of real IDs.
@@ 3.2 Acquisition Setup: Attacks
Fig. 1 shows visual examples of the physical attacks (print, screen and physical composite) available in our proposed FakeIDet2-db. Note that the personal information is redacted in the figure of the article to protect the privacy of the subject.
To manufacture the print attacks, we used a HP ScanJet 8270 scanner at 600 Dots Per Inch (DPI), creating a digital copy of the real ID. After that, we printed the fake IDs using an EPSON ET-2850 Wi-Fi printer and laminated them to improve the realism using a Fellowes Lunar A4 thermal laminator, providing print attacks with the characteristic glossy texture from real IDs. The acquisition process was the same as for the real IDs, yielding a total of 2,115 images for print attacks.
For screen attacks, we used a MacBook Pro 14” XDR screen, where the 2,115 images of real documents were displayed in full-screen and covering the whole smartphone camera preview. To increase variability across smartphone devices, each image was displayed and then captured using one randomly selected smartphone, ensuring that approximately one-third of the images were taken with each device.
Regarding the physical composite attacks, we cut sensitive sections from print attacks and placed them over the sensitive areas of real IDs. Regarding privacy considerations, we only cut the date of birth, the name and both surnames from the ID owner along with partial sections of the ID’s number. To provide even more variability, some physical composite attacks were created using cut-outs from several print attacks. The cut-outs of sensitive sections from a print attack were never displayed at the same time, further protecting the ID owner’s sensitive information. Additionally, we did not use the ID owner’s portrait image, as it is sensitive information which cannot be shared by any circumstances. After that, we captured composite attacks following the same procedure as for real IDs, described in Sec. 3.1, balancing the number of images of real and all types of attack across the database to 2,115.
Gathering all images from real and fake IDs, our database contains 8,460 total IDs images. From them, we selected 1,000 images, balancing them across different types of attacks and bona fide data (250 samples per class), different capture devices, and different illumination and height conditions. This sub-sampling aimed to avoid including images of real and fake IDs under all acquisition conditions, ensuring that fake ID detectors do not rely on superficial cues, such as illumination patterns, to distinguish between real and fake IDs, which could lead to overfitting [16].
@@ 3.3 Post-Processing: Anonymization and Patch Extraction
Fig. 4 provides graphical examples of the different anonymization configurations explored in the present study in order to facilitate the collaboration between ID Holders and AI Researchers in real-world fake ID detection scenarios. The criteria followed to remove sensitive information at each anonymization configuration is the following:
@@
    Pseudo-Anonymized: partial sections with sensitive information of real/fake IDs is displayed. This kind of anonymization is consistent along all images from a same ID, so no reconstruction is permitted.
@@
Referring to pseudo- and fully- anonymized configurations, we used EasyOCR [^8] https://github.com/JaidedAI/EasyOCR and GNU Image Manipulator Program (GIMP) to cover the sensitive regions with pitch black rectangles (with the color code (0,0,0) in the RGB spectrum). Since we had two levels of anonymization applied to the original, non-anonymized images, we obtained 2,000 images of real and fake IDs. Given that we had several images taken from the same ID, we took special care in covering the same partial sensitive sections across every image in the pseudo-anonymized configuration. Once all the IDs were anonymized, we processed them and extracted patches using PyTorch’s unfold method, specifying the same step and stride size as the patch size, to avoid overlapping. Patches with over 80% of pitch black content were discarded and the remaining ones were kept with a sampling probability of $p=0.9$, to further prevent reconstruction. After that, for each ID image, we created a folder where the extracted patches were stored using random numbers of six figures as the filename for patch (i.e., 762542.jpeg), so no spatial layout could be inferred.
An important point is the amount of patches available regarding anonymization and patch size configurations, which are depicted in Table 3 for real IDs and the different attacks considered. For instance, the Full-Anon, $128\times 128$ configuration shows a noticeable drop in terms of number of patches compared to the Non-Anon, $64\times 64$ configuration, having almost ten times less amount of patches. As we plan to release the FakeIDet2-db, the pseudo- and fully-anonymized ID configurations at $\times128$ and $\times64$ patch sizes will be available for privacy reasons, proving in the results section (Sec. 6) that these privacy-aware scenarios are very useful to facilitate collaborations between ID Holders and AI Researchers. If we consider the number of patches from said configurations, a total of 922,057 patches extracted from 2,000 images will be available for researchers to train and benchmark their proposed fake ID detectors.
@@
For completeness, Fig. 5 shows different examples of real/fake patches extracted from different IDs. One can see that the $\times128$ patch size contains much more information compared to their $\times64$ counterparts. Some small traces can be observed in the different attacks. For instance, the composite attacks might show the change of color and texture between the background and the altered section at the foreground. On the other hand, patches from the screen attacks might show subtle traces of the characteristic Moiré patterns.
4 Proposed Method: FakeIDet2
Our proposed fake ID detection method, FakeIDet2, proposes a privacy-aware, patch-level representation fusion. Fig. 6 provides a graphical representation of FakeIDet2. This comprises three main modules: the Privacy-Aware Patch Extractor, the Patch Embedding Extractor, and the Patch Embedding Fusion. The Privacy-Aware Patch Extractor module receives a complete ID, considering any anonymization level presented in Sec. 3.3. This first module divides the whole ID into patches, removing sensitive information if applicable (i.e., regions of the ID that are blacked out). After that, those patches are forwarded into the Patch Embedding Extractor, which uses a fine-tuned DINOv2 as a feature extractor to learn discriminative features at patch-level from each type of PAI, providing an embedding for each patch. Then, those embeddings are treated as tokens, and passed into a MHSA layer, which ponders the embeddings values depending on their importance in the sequence. After that, and attention pool layer selects discriminative features from the sequence, obtaining a single embedding, which is forwarded to an Multi-Layer Perceptron (MLP) that yields a final score per ID, indicating whether is real or fake. We describe next the technical details of the learnable modules of FakeIDet2.
Figure 6: Description of our proposed fake ID detector, FakeIDet2. Our proposal has three different stages: i) the whole ID is forwarded to the Privacy-Aware Patch Extractor, extracting patches at fixed size ($\times128$ or $\times64$); ii) those patches are forwarded into the Patch Embedding Extractor, which yields an embedding per patch; and iii) finally, those embeddings are used at the Patch Embedding Fusion module to generate a single embedding which is used by a MLP, providing a final score.

### v2

@@ 2.2 Fake ID Detection Methods and International Challenges
One of the first methods for the detection of fake IDs was presented in [24], where the authors trained a CNN by pixel classification using an internal database, reporting a 2.22% EER. Gonzalez and Tapia proposed in [15] a two-stage system which first used a neural network to evaluate if the ID is real or fake, evaluating digital PAIs of type composite and synthetic. The second network classified the ID between real or fake, evaluating physical PAIs such as print, display, and PVC. When concatenating both systems, the Bona-fide Presentation Classification Error Rate at a fixed decision threshold $\tau=0.01$ ($\textrm{BPCER}_{100}$) was 0.92%.
@@
A corroboration that the fake ID detection field is receiving increasing attention in recent years is the organization of international challenges at top conferences. The first challenge was introduced at IEEE IJCB 2024 [34]. The organizers did not provide data to prepare the competition, and the fake ID detectors submitted were evaluated on real IDs. The winner of the challenge achieved an EER of 21.87%. The second edition of this challenge has been hosted at IEEE IJCB 2025777https://sites.google.com/view/ijcb-pad-id-card-2025/, where the organizers proposed two tracks. The first track considered IDs created by the organizers using PVC cards as “real” documents, the rest as attacks, and the second track treated only official IDs as real. Only for the first track, a database was given to the participants. The winner of the first track reported an EER of 11.34%, while for the second track, the best team reported an EER of 6.36%, but the second and third teams reported EER s of 23.87% and 31.94%, respectively. These results show that this is a very challenging task, especially because of the lack of public real data. More recently, the first challenge including digital manipulations for fake ID detection (DeepID challenge) has been organized at IEEE/CVF ICCV 2025888https://deepid-iccv.github.io/. This challenge proposed two tracks: one for binary classification focused on PAD (i.e., detecting fake and real documents) and one for localization (i.e., providing a mask showing the tampered sections). The organizers provided a database for all participants, named FantasyID, which contained fake templates from different countries from different regions of the world. The “real” ID data was created by printing high quality fakes using PVC cards and taking images of them. After that, those images were digitally altered through face-swapping and text-inpainting methods. For the detection track, the evaluation was performed using the F1 score with a decision threshold fixed at $\tau=0.5$ on two datasets: the FantasyID evaluation set, which contained PAI s that were not available in the training set, and a private dataset provided by the company PXL Vision, with the same type of attacks as in the FantasyID dataset, but using official IDs as real. The final score, named Aggregated F1 score, was a pondered mean that gave more importance to the F1 score obtained over the private dataset, hence benefiting models that generalized better to out-of-distribution data. The winning team achieved an Aggregated F1 score of 0.8, obtaining an F1 score of 0.99 and 0.72 in both the FantasyID dataset and the private dataset, respectively. These results, although impressive, show that there is a present issue in domain adaptation between official real IDs and fake samples created under laboratory conditions that are considered “real”. Hence, providing privacy-aware mechanisms that allow sharing more realistic ID data without compromising individual’s personal information is fundamental to further advance in this research field.
2.3 State of the Art: Summary
@@
A clear trend can be observed on the basis of previous sections. Public databases in the literature do not have official real data available, and fake ID detectors trained using private databases with real data do not release the code or the weights to the research community, making it impossible to perform fair comparisons among state-of-the-art methods. Moreover, the characteristics of these databases are not always detailed (e.g., acquisition devices or number of unique IDs), which makes it even more difficult to advance in the field. The present study aims to advance through the following key contributions: i) the release of a new database, FakeIDet2-db, increasing the number of images from real and fake IDs, ii) the inclusion of a new kind of PAIs, named physical composite, to the proposed database, being the first one that contains this kind of attack, iii) a novel method for privacy-aware fake ID detection based on fusion of representations from individual patches, and iv) a new standard public benchmark that considers a wide variety of physical and synthetic attacks included in other public databases in the literature, considering IDs with different templates and demographic regions.
@@ 3 Proposed Database: FakeIDet2-db
This section describes our new FakeIDet2-db. Concretely, Sec. 3.1 and Sec. 3.2 provide all the details of the real and fake ID acquisition, whereas Sec. 3.3 describes the post-processing carried out for the different anonymization and patch size configurations.
@@ 3.1 Acquisition Setup: Real IDs
Figure 3: Main characteristics of our proposed public FakeIDet2-db. For the Fake IDs - Attacks, the personal information is redacted in this figure to protect the privacy of the subjects. Official Real IDs are based on template examples.
Fig. 3 summarizes the main characteristics of our proposed FakeIDet2-db. Regarding the number of unique IDs, our database comprises 47 different official Spanish IDs. The appearance of Spanish electronic IDs has changed over the years, which is a key aspect that has also been considered when capturing this database, including three different versions (see the Official Real IDs in Fig. 3). For the acquisition, we used three different smartphone models (Apple iPhone 15, Xiaomi Mi 9T Pro, and Redmi 9C NFC), to cover the range of sensor quality from high-end to low-end, respectively. Their main technical specifications are provided in Table 2 where we can see that both iPhone 15 and Mi 9T Pro have main camera sensors with 48MP while the Redmi 9C NFC only has 13MP. The iPhone camera is the best in terms of adaptability to low-light scenarios, given its aperture of ff/1.6, although the Mi 9T Pro is close with an aperture of ff/1.75. However, the Redmi 9C NFC struggles in these kind of low-light scenarios, since its aperture is greater, ff/$2.$ Regarding image resolution, the iPhone 15 is able to shoot ProRes images up to 8K resolution, but for our acquisition process, we set it to 4K, since both the Mi 9T Pro and the Redmi 9C NFC can only shoot images at that resolution. The photos of IDs were taken using a 4:3 aspect ratio. Three different heights were considered with respect to the document’s vertical: 10 cms, 12.5 cms, and 15 cms, on a 3D printed adjustable platform that we developed specifically for this research. This platform enabled acquiring IDs precisely, similar to [24], where they captured IDs using a mobile application that provided a template in the camera preview to match the ID borders.
To provide realistic variability, we also considered different light conditions in terms of room illumination: i) no-light, with the camera flash activated, ii) dim-light, with images taken with/without the camera flash activated, and iii) bright-light, with images taken with/without the camera flash activated.
Therefore, given the variability factors mentioned above, for a unique ID, we obtained 45 images (i.e., 3 smartphones $\times$ 3 heights $\times$ 5 light conditions) covering different conditions, generating a total of 2,115 (i.e., 47 $\times$ 45) images of real IDs.
@@ 3.2 Acquisition Setup: Attacks
Fig. 1 shows visual examples of physical attacks (print, screen, and physical composite) available in our proposed FakeIDet2-db. Note that personal information is redacted in the figure of the article to protect the privacy of the subject.
To manufacture the print attacks, we used an HP ScanJet 8270 scanner at 600 Dots Per Inch (DPI), creating a digital copy of the real ID. After that, we printed the fake IDs using an EPSON ET-2850 Wi-Fi printer and laminated them to improve the realism using a Fellowes Lunar A4 thermal laminator, providing print attacks with the characteristic glossy texture of real IDs. The acquisition process was the same as for real IDs, yielding a total of 2,115 images for print attacks.
For screen attacks, we used a MacBook Pro 14” XDR screen, where the 2,115 images of real documents were displayed in full-screen and covering the whole smartphone camera preview. To increase variability between smartphone devices, each image was displayed and then captured using a randomly selected smartphone, ensuring that approximately one third of the images were taken with each device.
Regarding physical composite attacks, we cut sensitive sections from print attacks and placed them in the sensitive areas of the real IDs. Regarding privacy considerations, we only cut the date of birth, the name, and both surnames from the ID owner along with partial sections of the ID’s number. To provide even more variability, some physical composite attacks were created using cut-outs of several print attacks. The cutouts of sensitive sections from a print attack were never displayed at the same time, further protecting the sensitive information of the ID owner. Additionally, we did not use the ID owner’s portrait image, as it is sensitive information that cannot be shared under any circumstances. After that, we captured composite attacks following the same procedure as for real IDs, described in Sec. 3.1, balancing the number of images of real and all types of attack across the database to 2,115.
Gathering all images from real and fake IDs, our database contains 8,460 total IDs images. From them, we selected 1,000 images, balancing them across different types of attacks and bona fide data (250 samples per class), different capture devices, and different illumination and height conditions. This subsampling aimed to avoid including images of real and fake IDs under all acquisition conditions, ensuring that fake ID detectors do not rely on superficial cues, such as illumination patterns, to distinguish between real and fake IDs, which could lead to overfitting.
@@ 3.3 Post-Processing: Anonymization and Patch Extraction
Fig. 4 provides graphical examples of the different anonymization configurations explored in the present study to facilitate collaboration between ID Holders and AI Researchers in real-world fake ID detection scenarios. The criteria followed to remove sensitive information at each anonymization configuration are the following:
@@
    Pseudo-Anonymized: partial sections with sensitive information of real/fake IDs is displayed. This kind of anonymization is consistent along all images from the same ID, so no reconstruction is allowed.
@@
Referring to pseudo- and fully-anonymized configurations, we used EasyOCR [^9] https://github.com/JaidedAI/EasyOCR and the GNU Image Manipulator Program (GIMP) to cover the sensitive regions with pitch black rectangles (with the color code (0,0,0) in the RGB spectrum). Since we had two levels of anonymization applied to the original, non-anonymized images, we obtained 2,000 images of real and fake IDs. Given that we had several images taken from the same ID, we took special care to cover the same partial sensitive sections across all images in the pseudo-anonymized configuration. Once all IDs were anonymized, we processed them and extracted patches using PyTorch’s unfold method, specifying the same step and stride size as the patch size to avoid overlapping. Patches with a pitch black content greater than 80% were discarded and the remaining were kept with a sampling probability of $p=0.9$ to further prevent reconstruction. After that, for each ID image, we created a folder where the extracted patches were stored using random numbers of six figures as the filename of the patch (e.g., 762542.jpeg), so no spatial layout could be inferred.
An important point is the number of patches available with respect to anonymization and patch size configurations, which are depicted in Table 3 for real IDs and the different attacks considered. For example, the Full-Anon, $128\times 128$ configuration shows a noticeable drop in terms of the number of patches compared to the Non-Anon, $64\times 64$ configuration, which has almost ten times fewer patches. As we plan to release FakeIDet2-db, the pseudo- and fully-anonymized ID configurations at $\times128$ and $\times64$ patch sizes will be available for privacy reasons, as proved in the results section (Sec. 6) that these privacy-aware scenarios are very useful to facilitate collaborations between ID Holders and AI Researchers. If we consider the number of patches from the said configurations, a total of 922,057 patches extracted from 2,000 images will be available to researchers to train and benchmark their proposed fake ID detectors.
@@
For completeness, Fig. 5 shows different examples of real/fake patches extracted from different IDs. One can see that the $\times128$ patch size contains much more information compared to their $\times64$ counterparts. Some small traces can be observed in the different attacks. For example, composite attacks might show the change in color and texture between the background and the altered section in the foreground. However, patches from screen attacks might show subtle traces of the characteristic Moiré patterns.
4 Proposed Method for Fake ID Detection: FakeIDet2
Our proposed fake ID detection method, FakeIDet2, proposes a privacy-aware, patch-level representation fusion [14]. Fig. 6 provides a graphical representation of FakeIDet2. This comprises three main modules: the Privacy-Aware Patch Extractor, the Patch Embedding Extractor, and the Patch Embedding Fusion. The Privacy-Aware Patch Extractor module receives a complete ID, considering any anonymization level presented in Sec. 3.3. This first module divides the entire ID into patches, removing sensitive information if applicable (i.e., the regions of the ID that are blacked out). After that, those patches are forwarded to the Patch Embedding Extractor, which uses a fine-tuned DINOv2 as a feature extractor to learn discriminative features at the patch level of each type of PAI, providing an embedding for each patch. Then, those embeddings are treated as tokens, and passed into a MHSA layer, which ponders the embeddings values depending on their importance in the sequence. After that, an attention pool layer selects discriminative features from the sequence, obtaining a single embedding, which is forwarded to a Multi-Layer Perceptron (MLP) that yields a final score per ID, indicating whether it is real or fake. Next, we describe the technical details of the learnable modules of FakeIDet2.
Figure 6: General diagram of our proposed fake ID detector, FakeIDet2. Our proposal has three different stages: i) the entire ID is forwarded to the Privacy-Aware Patch Extractor, extracting patches of fixed size ($\times128$ or $\times64$); ii) those patches are forwarded into the Patch Embedding Extractor, which yields an embedding per patch; and iii) finally, those embeddings are used at the Patch Embedding Fusion module to generate a single embedding which is used by an MLP, providing a final score that is thresholded to obtain the final decision: real or fake ID.